# Dozerformer: Sequence Adaptive Sparse Transformer for Multivariate Time Series Forecasting

### v1

@@ 4.4.1 Local, Stride, and Vary Components Ablation
To investigate the impact of the Local, Stride, and Vary components individually, we conducted experiments by using only one of these components in the Dozer Attention. The results are summarized in Table 3, focusing on forecasting horizons of 96, 192, 336, and 720. The hyper-parameter of each component follows the setting of the main results. Overall, each component working alone performed slightly worse than the Dozer Attention, showing the effectiveness of each component in capturing the essential attributes of MTS data.
@@ 4.4.2 Attention Mechanism Comparison
To demonstrate the effectiveness of the proposed Dozer attention mechanism, we conducted a comparative analysis by replacing it with other attention mechanisms in the Dozerformer framework. Specifically, we compared it with canonical full attention Vaswani et al. (2017), Auto-Correlation Wu et al. (2021), Frequency Enhanced Auto-Correlation Zhou et al. (2022), and ProbSparse attention Zhou et al. (2021). The results of this comparison are presented in Table 4. The Dozer Attention outperformed other attention mechanisms, securing the top position with 22 best-case scenarios out of 32. Canonical full attention performed as the second-best, achieving the best results in 9 cases. Significantly, the Dozer Attention reduced the average number of query-key pairs to 31.7% and 28.4% for self-attention in the encoder and cross-attention in the decoder, respectively, compared to full attention. This reduction represents a substantial decrease in computational complexity by eliminating the majority of query-key pairs. The Auto-Correlation performed notably worse than the proposed Dozer Attention, with its MSE and MAE being 20.1% and 13.6% higher, respectively. The Frequency Enhanced Auto-Correlation exhibited performance that was 13.8% worse in MSE and 8% worse in MAE compared to the Dozer Attention. The ProbSparse Attention, while still competitive, performed slightly worse, with 5.7% higher MSE and 2% higher MAE than the Dozer Attention. We conclude that the proposed Dozer attention is just as capable as the canonical full attention but significantly more efficient and powerful than recent state-of-the-art attention mechanisms.
@@ 4.5.1 Effect of local, stride, and vary size
To investigate the effect of local size $w$, stride size $s$, and vary size $v$ to MTS forecasting accuracy, we conduct experiments and present the results in Figure 6. From the computational complexity aspect, smaller $w$, $s$, and $v$ indicate less computation. By increasing $w$, $s$, and $v$, the Dozer attention gets less sparse. However, including more query-key pairs doesn’t improve the forecasting accuracy. Indicating the importance of removing less correlated historical time steps.
@@ 4.5.2 Effect of look-back window size
The size of the look-back window, denoted as $I$, impacts primarily the stride component of the Dozerformer. Figure 7 illustrates the effect of look-back window size on seven methods at horizons 96 and 720. For the 96-hour horizon, no method benefits from a larger look-back window size except for the Exchange-Rate dataset. The Dozerformer consistently outperforms all baseline methods, indicating that time steps distant from the prediction target are less correlated and can even degrade forecasting accuracy. However, for the 720-hour horizon, a longer look-back window size enhances the Dozerformer’s forecasting accuracy. The Exchange-Rate dataset, known for its locality, performs optimally with a small look-back window size for all methods. In conclusion, the optimal input sequence length varies with the forecasting horizon; shorter horizons favor a shorter look-back window, and vice versa.
@@ 5 Conclusions
This paper introduced a sequence adaptive sparse transformer framework named Dozerformer for MTS forecasting. The Dozerformer incorporates a sparse Dozer attention mechanism which consists of local, stride, and vary components to selectively attend each query to keys based on the characteristic of MTS data and forecasting horizons. We demonstrated the superior performance of Dozerformer against recent state-of-the-art methods in both accuracy and efficiency. For future work, we plan to further modify the Dozer attention to model multi-scale temporal dependencies.
@@
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
@@ Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
@@ A.1 Datasets Details
  • Traffic [^2] http://pems.dot.ca.gov: This dataset sourced from the California Department of Transportation, comprises road occupancy rate measurements, ranging from 0 to 1, gathered from 862 freeway sites within the San Francisco Bay area. This publicly available data spans over a decade. Lai lai2018modeling meticulously collected hourly data over a period of 48 months (2015-2016), resulting in a total of 17,544 data samples.
@@
  • Weather [^4] https://www.bgc-jena.mpg.de/wetter/: This dataset encompasses 21 meteorological indicators, with recordings captured at 10-minute intervals over the span of a year in Germany, resulting in a total of 52,696 samples.
@@ A.2 Detailed Computation of the Dozerformer
The transformer encoder-decoder pair utilizing the Dozer attention is employed as the seasonal model. The DI embedding Zhang et al. (2023) is utilized to embed the raw MTS into feature maps and partition them into patches.
@@
A.3 More Experimental Results
A.3.1 The quantitative results in MASE
@@
A.3.2 The Model complexity of all methods
To assess the efficiency of the proposed Dozerformer in comparison with baseline methods, we selected ETTh1 as an illustrative example and evaluated the Params (total number of learnable parameters), FLOPs (floating-point operations), and Memory (maximum GPU memory consumption). With a historical length set at 720 and a forecasting horizon of 96, Table 7 presents the quantitative results of this efficiency comparison. Dlinear Zeng et al. (2023), characterized by a straightforward architecture employing a linear layer to directly generate forecasting values from historical records, achieved the best efficiency among the methods considered. Notably, the proposed Dozerformer exhibited superior efficiency compared to other baseline methods based on CNN or transformers. It is essential to highlight that patchTST Nie et al. (2023) is the only method that did not significantly lag behind Dozerformer in accuracy. However, its parameter count is six times greater than that of Dozerformer, and FLOPs are 17 times larger.
Table 7: The Model complexity results in validation dataset for input length of 720 and forecasting horizon 96
Params (M) 1.61 10.73 0.13 52.99 18.78 42.36 215.01 9.05 11.32
A.3.3 Visualization of forecasting results
@@
A.3.4 More Ablation Studies
Figure 6 presents the results of an ablation study examining the impact of forecasting accuracy under different settings of local sizes, stride sizes, and vary starting sizes—key hyperparameters of the core Local, Stride, and Vary components of the Dozerformer. These hyperparameters influence the sparsity of the attention matrix, consequently affecting the efficiency of the Dozerformer in terms of computation and memory usage. A smaller local size, vary starting size, and a larger stride value result in a more sparse attention matrix, contributing to increased efficiency. It is noteworthy that an increase in historical time steps utilized in the attention mechanism does not necessarily lead to enhanced accuracy. The optimal local size varies based on the dataset’s characteristics, but we observe higher MSE values as the local size increases and more time steps are utilized. Regarding the stride value, it reaches an optimum of 9 for the ETTh1 and ETTm1 datasets, while it is 1 for the Weather datasets. For the vary value, a small value of 1 produces good performance in terms of accuracy and efficiency. In summary, we conclude that the selection of these hyperparameters varies across datasets. However, a sparse matrix, in general, performs better than a more dense attention matrix.
The historical time step size is a crucial parameter for time series forecasting tasks, as it determines the model’s receptive field. In theory, a longer sequence length should enhance forecasting accuracy by providing the model with more historical information. However, most methods encounter challenges in aligning with this assumption. Figure 7 illustrates how different methods are influenced by the input sequence length. Remarkably, the Dozerformer consistently outperforms other methods for various sequence lengths in the ETTh1, ETTm1, and Weather datasets. The Exchange-Rate dataset is an exception; due to its lack of seasonality, all methods experienced a decline in accuracy with longer input sequences. Notably, an increase in input sequence length does not necessarily lead to improved accuracy for the Dozerformer. In contrast, for the remaining baseline methods, their performance tends to deteriorate as the input sequence length increases. We attribute this trend to longer input sequences incorporating more historical time steps that lack correlation with the forecasting target, thereby degrading the performance of these methods. The Dozerformer, by nature of its sparsity, is less affected, as it naturally excludes uncorrelated time steps from its attention mechanism.

### v2

@@ 4.3 Computational Efficiency
We selected ETTh1 as an illustrative example and evaluated the Params (total number of learnable parameters), FLOPs (floating-point operations), and Memory (maximum GPU memory consumption). With a historical length set at 720 and a forecasting horizon of 96, Table 3 presents the quantitative results of this efficiency comparison. Dlinear Zeng et al. (2023), characterized by a straightforward architecture employing a linear layer to directly generate forecasting values from historical records, achieved the best efficiency among the methods considered. Notably, the proposed Dozerformer exhibited superior efficiency compared to other baseline methods based on CNN or transformers. It is essential to highlight that patchTST Nie et al. (2023) is the only method that did not significantly lag behind Dozerformer in accuracy. However, its parameter count is six times greater than that of Dozerformer, and FLOPs are 17 times larger.
Table 3: The Model complexity results in ETTh1 dataset for input length of 720 and forecasting horizon 96
Params (M) 1.61 10.73 0.13 52.99 18.78 42.36 215.01 9.05 11.32
@@ 4.4.1 Local, Stride, and Vary Components Ablation
To investigate the impact of the Local, Stride, and Vary components individually, we conducted experiments by using only one of these components in the Dozer Attention. The results are summarized in Table 4, the mean value across four horizons (96, 192, 336, and 720) for seed 1 was presented. The hyper-parameter of each component follows the setting of the main results. Overall, each component working alone performed slightly worse than the Dozer Attention, showing the effectiveness of each component in capturing the essential attributes of MTS data.
@@ 4.4.2 Attention Mechanism Comparison
To demonstrate the effectiveness of the proposed Dozer attention mechanism, we conducted a comparative analysis by replacing it with other attention mechanisms in the Dozerformer framework. Specifically, we compared it with canonical full attention Vaswani et al. (2017), Auto-Correlation Wu et al. (2021), Frequency Enhanced Auto-Correlation Zhou et al. (2022), and ProbSparse attention Zhou et al. (2021). The results of this comparison are presented in Table 5. The Dozer Attention outperformed other attention mechanisms, securing the top position with 22 best-case scenarios out of 32. Canonical full attention performed as the second-best, achieving the best results in 9 cases. Significantly, the Dozer Attention reduced the average number of query-key pairs to 31.7% and 28.4% for self-attention in the encoder and cross-attention in the decoder, respectively, compared to full attention. This reduction represents a substantial decrease in computational complexity by eliminating the majority of query-key pairs. The Auto-Correlation performed notably worse than the proposed Dozer Attention, with its MSE and MAE being 20.1% and 13.6% higher, respectively. The Frequency Enhanced Auto-Correlation exhibited performance that was 13.8% worse in MSE and 8% worse in MAE compared to the Dozer Attention. The ProbSparse Attention, while still competitive, performed slightly worse, with 5.7% higher MSE and 2% higher MAE than the Dozer Attention. We conclude that the proposed Dozer attention is just as capable as the canonical full attention but significantly more efficient than recent state-of-the-art attention mechanisms.
@@ 4.5.1 Effect of local, stride, and vary size
To investigate the effect of local size $w$, stride size $s$, and vary size $v$ to MTS forecasting accuracy, we conduct experiments and present the results in Figure 4. These hyperparameters influence the sparsity of the attention matrix, consequently affecting the efficiency of the Dozerformer in terms of computation and memory usage. A smaller local size $w$, vary starting size $v$, and a larger stride size $s$ result in a more sparse attention matrix, contributing to increased efficiency. It is noteworthy that an increase in historical time steps utilized in the attention mechanism does not necessarily lead to enhanced accuracy. The optimal local size varies based on the dataset’s characteristics, but we observe higher MSE values as the local size increases and more time steps are utilized. Regarding the stride size, it reaches an optimum of 9 for the ETTh1 and ETTm1 datasets, while it is 1 for the Weather datasets. For the vary size, a small value of 1 produces good performance in terms of accuracy and efficiency. In summary, we conclude that the selection of these hyperparameters varies across datasets. However, a sparse matrix, in general, performs better than a more dense attention matrix.
@@ 4.5.2 Effect of look-back window size
The size of the look-back window, denoted as $I$, impacts primarily the stride component of the Dozerformer, as it determines the model’s receptive field. Figure 5 illustrates the effect of look-back window size on seven methods at horizons 96 and 720. Remarkably, the Dozerformer consistently outperforms other methods for various sequence lengths in the ETTh1, ETTm1, and Weather datasets. The Exchange-Rate dataset is an exception; due to its lack of seasonality, all methods experienced a decline in accuracy with longer input sequences. Notably, an increase in input sequence length does not necessarily lead to improved accuracy for the Dozerformer. In contrast, for the remaining baseline methods, their performance tends to deteriorate as the input sequence length increases. We attribute this trend to longer input sequences incorporating more historical time steps that lack correlation with the forecasting target, thereby degrading the performance of these methods. The Dozerformer, by nature of its sparsity, is less affected, as it naturally excludes uncorrelated time steps from its attention mechanism.
@@ 5 Conclusions
This paper introduced a sequence adaptive sparse transformer framework named Dozerformer for MTS forecasting. The Dozerformer incorporates a sparse Dozer attention mechanism which consists of local, stride, and vary components to selectively attend each query to keys based on the characteristic of MTS data and forecasting horizons. We demonstrated the superior performance of Dozerformer against recent state-of-the-art methods in both accuracy and efficiency. The design of the Local, Stride, and Vary components requires proper selection of their sizes, necessitating an understanding of the dataset characteristics to optimally set these hyperparameters. For future work, we plan to further modify the Dozer attention to model multi-scale temporal dependencies and adaptively select subsets of historical time steps for the sparse attention mechanism.
@@
YZ and RW developed the approach. YZ implemented the method in Pytorch and conducted experiments. YZ wrote the manuscript. RW, SD, and FH proofread the manuscript. All authors reviewed the manuscript.
@@ Acknowledgments
This material is based in part upon work supported by: The National Science Foundation under grant number(s) NSF awards 2142428, 2142360, OIA-2019609, and OIA-2148788. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Datasets Details
  • Traffic [^2] http://pems.dot.ca.gov: This dataset sourced from the California Department of Transportation, comprises road occupancy rate measurements, ranging from 0 to 1, gathered from 862 freeway sites within the San Francisco Bay area. This publicly available data spans over a decade. Lai Lai et al. (2018a) meticulously collected hourly data over 48 months (2015-2016), resulting in a total of 17,544 data samples.
@@
  • Weather [^4] https://www.bgc-jena.mpg.de/wetter/: This dataset encompasses 21 meteorological indicators, with recordings captured at 10-minute intervals over a year in Germany, resulting in a total of 52,696 samples.
@@
The datasets were divided into distinct subsets: training, validation, and test sets. Specifically, ETTh1, ETTh2, ETTm1, and ETTm2 were divided in ratios of 0.6, 0.2, and 0.2, respectively. The other five datasets were allocated with a division ratio of 0.7, 0.1, and 0.2.
@@ A.2 Detailed Computation of the Dozerformer
The transformer encoder-decoder pair utilizing the Dozer attention is employed as the seasonal model. The DI embedding Zhang et al. (2024a) is utilized to embed the raw MTS into feature maps and partition them into patches.
@@
Appendix B More Experimental Results
B.1 The quantitative results in MASE
@@
B.2 Visualization of forecasting results